What amount must Harold save each month for the next 4 years if he wants to accumulate $40,000 and he can earn an annual rate of 9 percent?  His first deposit will be one month from today. (Choose the closest answer)
Answer #1

Future value = pmtx ((1+r)^n-1))/ PMT = FV = ((1+r)^n-1)/r Save each month = PMT = $695.40 ( 140,000/(((1.0075-48)-1)/0.0075)